OpenStudy (anonymous):

Can somebody help me? I would Also like to see the steps if possible. What is the axis of symmetry of the graph of y = –3(x + 8)2 + 5? Answer x = 8 x = –3 x = –8 x = 5

OpenStudy (anonymous):

y=-3(x+8)2 + 5 y=(-3x -24)2+5 y=-6x-48+5 y=-6x-43 try drawing the graph using this equation and find the axis of sym